# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ Appendix I Machine Learning Efficiency Models
acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.959_{\pm 0.012}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
beijing $0.712_{\pm 0.001}$ $0.739_{\pm 0.002}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.788_{\pm 0.011}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
lending $0.030_{\pm 0.000}$ $0.042_{\pm 0.001}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.268_{\pm 0.004}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
news $1.001_{\pm 0.002}$ $1.180_{\pm 0.107}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.374_{\pm 0.028}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
nmes $1.001_{\pm 0.003}$ $1.112_{\pm 0.044}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $0.535_{\pm 0.013}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
Appendix S Ablation Study Details
@@
acsincome $0.131_{\pm 0.003}$ $0.119_{\pm 0.004}$ $0.124_{\pm 0.006}$ $0.129_{\pm 0.004}$ $0.129_{\pm 0.003}$
adult $0.131_{\pm 0.007}$ $0.128_{\pm 0.008}$ $0.168_{\pm 0.017}$ $0.107_{\pm 0.011}$ $0.125_{\pm 0.009}$
beijing $0.065_{\pm 0.009}$ $0.066_{\pm 0.012}$ $0.067_{\pm 0.011}$ $0.067_{\pm 0.010}$ $0.073_{\pm 0.009}$
churn $0.244_{\pm 0.015}$ $0.272_{\pm 0.034}$ $0.299_{\pm 0.066}$ $0.264_{\pm 0.012}$ $0.289_{\pm 0.043}$
@@
acsincome $0.025_{\pm 0.000}$ $0.025_{\pm 0.001}$ $0.025_{\pm 0.001}$ $0.024_{\pm 0.001}$ $0.024_{\pm 0.000}$
adult $0.012_{\pm 0.001}$ $0.013_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.014_{\pm 0.001}$ $0.013_{\pm 0.001}$
beijing $0.004_{\pm 0.001}$ $0.006_{\pm 0.002}$ $0.005_{\pm 0.003}$ $0.004_{\pm 0.002}$ $0.005_{\pm 0.002}$
churn $0.010_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.009_{\pm 0.004}$ $0.010_{\pm 0.002}$ $0.012_{\pm 0.002}$
@@
acsincome $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
adult $0.004_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$
beijing $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
churn $0.006_{\pm 0.001}$ $0.006_{\pm 0.000}$ $0.006_{\pm 0.001}$ $0.006_{\pm 0.001}$ $0.007_{\pm 0.001}$
@@
acsincome $0.534_{\pm 0.002}$ $0.534_{\pm 0.001}$ $0.538_{\pm 0.003}$ $0.532_{\pm 0.002}$ $0.532_{\pm 0.004}$
adult $0.597_{\pm 0.002}$ $0.593_{\pm 0.001}$ $0.615_{\pm 0.003}$ $0.580_{\pm 0.003}$ $0.588_{\pm 0.002}$
beijing $0.953_{\pm 0.002}$ $0.959_{\pm 0.001}$ $0.952_{\pm 0.003}$ $0.953_{\pm 0.001}$ $0.949_{\pm 0.001}$
churn $0.557_{\pm 0.014}$ $0.573_{\pm 0.014}$ $0.564_{\pm 0.012}$ $0.541_{\pm 0.015}$ $0.533_{\pm 0.007}$
@@
acsincome $7.673_{\pm 0.017}$ $8.335_{\pm 0.064}$ $8.222_{\pm 0.035}$ $8.305_{\pm 0.021}$ $8.352_{\pm 0.025}$ $8.322_{\pm 0.047}$
adult $1.870_{\pm 0.000}$ $1.221_{\pm 0.018}$ $1.294_{\pm 0.015}$ $1.252_{\pm 0.014}$ $1.427_{\pm 0.008}$ $1.231_{\pm 0.011}$
beijing $0.385_{\pm 0.000}$ $0.545_{\pm 0.001}$ $0.559_{\pm 0.002}$ $0.539_{\pm 0.003}$ $0.541_{\pm 0.002}$ $0.489_{\pm 0.001}$
churn $0.347_{\pm 0.000}$ $0.307_{\pm 0.016}$ $0.326_{\pm 0.009}$ $0.294_{\pm 0.022}$ $0.298_{\pm 0.013}$ $0.274_{\pm 0.021}$
@@
adult $0.797_{\pm 0.000}$ $0.788_{\pm 0.001}$ $0.788_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.788_{\pm 0.002}$ $0.787_{\pm 0.001}$
churn $0.873_{\pm 0.003}$ $0.856_{\pm 0.008}$ $0.856_{\pm 0.014}$ $0.857_{\pm 0.007}$ $0.849_{\pm 0.006}$ $0.852_{\pm 0.006}$
@@
adult $0.915_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.910_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$
churn $0.964_{\pm 0.001}$ $0.962_{\pm 0.002}$ $0.961_{\pm 0.003}$ $0.960_{\pm 0.002}$ $0.961_{\pm 0.001}$ $0.962_{\pm 0.001}$
@@
acsincome $0.804_{\pm 0.012}$ $0.815_{\pm 0.009}$ $0.813_{\pm 0.018}$ $0.823_{\pm 0.017}$ $0.814_{\pm 0.014}$ $0.811_{\pm 0.014}$
beijing $0.712_{\pm 0.001}$ $0.782_{\pm 0.004}$ $0.785_{\pm 0.004}$ $0.778_{\pm 0.005}$ $0.776_{\pm 0.004}$ $0.770_{\pm 0.005}$
Appendix T Training and Sampling Times Details
@@
Table 29: Sample times in seconds per 1000 samples.

### v4

@@ Appendix I Tuning of the Detection Model
acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.955_{\pm 0.010}$ $0.827_{\pm 0.009}$ $0.807_{\pm 0.008}$ $0.812_{\pm 0.004}$
beijing $0.711_{\pm 0.001}$ $0.742_{\pm 0.002}$ $0.779_{\pm 0.007}$ $1.050_{\pm 0.010}$ $1.295_{\pm 0.016}$ $0.799_{\pm 0.007}$ $0.849_{\pm 0.004}$ $0.789_{\pm 0.010}$ $0.772_{\pm 0.003}$ $0.766_{\pm 0.003}$ $0.762_{\pm 0.002}$
lending $0.030_{\pm 0.000}$ $0.042_{\pm 0.001}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.789_{\pm 0.033}$ $0.305_{\pm 0.006}$ $0.071_{\pm 0.001}$ $0.066_{\pm 0.002}$ $0.062_{\pm 0.002}$
news $1.001_{\pm 0.002}$ $1.180_{\pm 0.107}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.171_{\pm 0.006}$ $1.302_{\pm 0.074}$ $0.397_{\pm 0.037}$ $0.848_{\pm 0.081}$ $0.752_{\pm 0.067}$ $0.717_{\pm 0.045}$
nmes $1.001_{\pm 0.003}$ $1.112_{\pm 0.044}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.200_{\pm 0.054}$ $1.137_{\pm 0.052}$ $0.563_{\pm 0.005}$ $1.154_{\pm 0.052}$ $1.109_{\pm 0.055}$ $1.136_{\pm 0.074}$
Appendix T Ablation Study Details
@@
acsincome $0.138_{\pm 0.004}$ $0.138_{\pm 0.003}$ $0.135_{\pm 0.001}$ $0.152_{\pm 0.001}$ $0.135_{\pm 0.002}$
adult $0.135_{\pm 0.006}$ $0.113_{\pm 0.004}$ $0.159_{\pm 0.015}$ $0.100_{\pm 0.010}$ $0.125_{\pm 0.005}$
bank $0.478_{\pm 0.019}$ $0.243_{\pm 0.015}$ $0.269_{\pm 0.012}$ $0.196_{\pm 0.012}$ $0.231_{\pm 0.009}$
beijing $0.076_{\pm 0.009}$ $0.070_{\pm 0.007}$ $0.067_{\pm 0.004}$ $0.068_{\pm 0.006}$ $0.073_{\pm 0.007}$
churn $0.294_{\pm 0.053}$ $0.280_{\pm 0.043}$ $0.262_{\pm 0.030}$ $0.239_{\pm 0.038}$ $0.255_{\pm 0.021}$
covertype $1.178_{\pm 0.193}$ $2.099_{\pm 0.010}$ $1.974_{\pm 0.136}$ $1.808_{\pm 0.010}$ $1.357_{\pm 0.155}$
default $0.905_{\pm 0.110}$ $0.799_{\pm 0.123}$ $0.727_{\pm 0.132}$ $0.521_{\pm 0.124}$ $0.641_{\pm 0.130}$
diabetes $0.719_{\pm 0.049}$ $1.435_{\pm 0.021}$ $1.397_{\pm 0.005}$ $1.230_{\pm 0.061}$ $1.213_{\pm 0.029}$
lending $1.480_{\pm 0.046}$ $1.127_{\pm 0.083}$ $1.178_{\pm 0.059}$ $1.295_{\pm 0.044}$ $1.239_{\pm 0.090}$
news $2.484_{\pm 0.138}$ $2.136_{\pm 0.417}$ $1.973_{\pm 0.417}$ $2.016_{\pm 0.483}$ $1.811_{\pm 0.295}$
nmes $0.483_{\pm 0.048}$ $0.421_{\pm 0.032}$ $0.457_{\pm 0.041}$ $0.450_{\pm 0.041}$ $0.444_{\pm 0.075}$
@@
acsincome $0.016_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.026_{\pm 0.001}$ $0.032_{\pm 0.001}$ $0.022_{\pm 0.001}$
adult $0.010_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.013_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.015_{\pm 0.001}$
bank $0.008_{\pm 0.000}$ $0.017_{\pm 0.000}$ $0.011_{\pm 0.001}$ $0.011_{\pm 0.001}$ $0.011_{\pm 0.001}$
beijing $0.005_{\pm 0.001}$ $0.005_{\pm 0.004}$ $0.004_{\pm 0.002}$ $0.004_{\pm 0.001}$ $0.006_{\pm 0.001}$
churn $0.015_{\pm 0.003}$ $0.011_{\pm 0.002}$ $0.009_{\pm 0.003}$ $0.010_{\pm 0.003}$ $0.010_{\pm 0.002}$
covertype $0.002_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$
default $0.014_{\pm 0.001}$ $0.015_{\pm 0.001}$ $0.012_{\pm 0.002}$ $0.017_{\pm 0.002}$ $0.016_{\pm 0.001}$
diabetes $0.024_{\pm 0.000}$ $0.030_{\pm 0.000}$ $0.022_{\pm 0.000}$ $0.024_{\pm 0.001}$ $0.025_{\pm 0.000}$
lending $0.055_{\pm 0.001}$ $0.055_{\pm 0.002}$ $0.055_{\pm 0.001}$ $0.057_{\pm 0.001}$ $0.057_{\pm 0.001}$
news $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.003_{\pm 0.001}$ $0.004_{\pm 0.001}$ $0.003_{\pm 0.001}$
nmes $0.008_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.008_{\pm 0.002}$ $0.009_{\pm 0.001}$ $0.007_{\pm 0.002}$
@@
acsincome $0.005_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.001_{\pm 0.000}$ $0.001_{\pm 0.000}$
adult $0.008_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.003_{\pm 0.000}$
bank $0.007_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.005_{\pm 0.001}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.001}$
beijing $0.006_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$ $0.002_{\pm 0.000}$
churn $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.007_{\pm 0.002}$ $0.006_{\pm 0.001}$
covertype $0.003_{\pm 0.000}$ $0.020_{\pm 0.000}$ $0.019_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.009_{\pm 0.000}$
default $0.003_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.004_{\pm 0.000}$ $0.003_{\pm 0.000}$ $0.003_{\pm 0.000}$
diabetes $0.021_{\pm 0.000}$ $0.042_{\pm 0.000}$ $0.041_{\pm 0.000}$ $0.033_{\pm 0.000}$ $0.034_{\pm 0.000}$
lending $0.012_{\pm 0.000}$ $0.011_{\pm 0.000}$ $0.012_{\pm 0.000}$ $0.013_{\pm 0.000}$ $0.013_{\pm 0.000}$
news $0.009_{\pm 0.000}$ $0.006_{\pm 0.000}$ $0.007_{\pm 0.000}$ $0.005_{\pm 0.000}$ $0.006_{\pm 0.000}$
nmes $0.008_{\pm 0.001}$ $0.007_{\pm 0.001}$ $0.008_{\pm 0.000}$ $0.008_{\pm 0.001}$ $0.006_{\pm 0.000}$
@@
acsincome $0.547_{\pm 0.001}$ $0.533_{\pm 0.003}$ $0.540_{\pm 0.002}$ $0.546_{\pm 0.003}$ $0.527_{\pm 0.002}$
adult $0.640_{\pm 0.003}$ $0.595_{\pm 0.002}$ $0.621_{\pm 0.002}$ $0.581_{\pm 0.002}$ $0.590_{\pm 0.004}$
bank $0.880_{\pm 0.001}$ $0.739_{\pm 0.006}$ $0.798_{\pm 0.001}$ $0.662_{\pm 0.005}$ $0.701_{\pm 0.005}$
beijing $0.699_{\pm 0.003}$ $0.658_{\pm 0.004}$ $0.615_{\pm 0.002}$ $0.626_{\pm 0.002}$ $0.620_{\pm 0.005}$
churn $0.710_{\pm 0.008}$ $0.610_{\pm 0.006}$ $0.580_{\pm 0.011}$ $0.560_{\pm 0.019}$ $0.541_{\pm 0.008}$
covertype $0.887_{\pm 0.002}$ $0.991_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.984_{\pm 0.001}$ $0.981_{\pm 0.000}$
default $0.925_{\pm 0.002}$ $0.816_{\pm 0.003}$ $0.774_{\pm 0.003}$ $0.759_{\pm 0.004}$ $0.793_{\pm 0.005}$
diabetes $0.762_{\pm 0.001}$ $0.895_{\pm 0.001}$ $0.870_{\pm 0.002}$ $0.851_{\pm 0.001}$ $0.837_{\pm 0.001}$
lending $0.989_{\pm 0.002}$ $0.938_{\pm 0.011}$ $0.958_{\pm 0.003}$ $0.944_{\pm 0.005}$ $0.955_{\pm 0.003}$
news $0.993_{\pm 0.001}$ $0.956_{\pm 0.002}$ $0.965_{\pm 0.002}$ $0.946_{\pm 0.002}$ $0.950_{\pm 0.002}$
nmes $0.663_{\pm 0.014}$ $0.651_{\pm 0.007}$ $0.667_{\pm 0.009}$ $0.648_{\pm 0.015}$ $0.633_{\pm 0.010}$
@@
acsincome $7.673_{\pm 0.017}$ $8.550_{\pm 0.055}$ $8.342_{\pm 0.037}$ $8.284_{\pm 0.030}$ $8.311_{\pm 0.041}$ $8.397_{\pm 0.062}$
adult $1.870_{\pm 0.000}$ $1.231_{\pm 0.016}$ $1.318_{\pm 0.009}$ $1.296_{\pm 0.012}$ $1.520_{\pm 0.011}$ $1.327_{\pm 0.015}$
bank $2.369_{\pm 0.000}$ $1.583_{\pm 0.006}$ $2.016_{\pm 0.007}$ $2.069_{\pm 0.014}$ $2.187_{\pm 0.009}$ $1.913_{\pm 0.006}$
beijing $0.385_{\pm 0.000}$ $0.592_{\pm 0.001}$ $0.537_{\pm 0.002}$ $0.505_{\pm 0.001}$ $0.511_{\pm 0.001}$ $0.473_{\pm 0.002}$
churn $0.347_{\pm 0.000}$ $0.526_{\pm 0.014}$ $0.383_{\pm 0.014}$ $0.327_{\pm 0.022}$ $0.333_{\pm 0.020}$ $0.278_{\pm 0.012}$
covertype $0.529_{\pm 0.001}$ $0.871_{\pm 0.009}$ $1.919_{\pm 0.016}$ $1.872_{\pm 0.022}$ $1.698_{\pm 0.024}$ $1.594_{\pm 0.009}$
default $1.812_{\pm 0.000}$ $1.321_{\pm 0.015}$ $1.333_{\pm 0.021}$ $1.306_{\pm 0.020}$ $1.484_{\pm 0.012}$ $1.355_{\pm 0.017}$
diabetes $15.608_{\pm 0.055}$ $13.299_{\pm 0.020}$ $14.958_{\pm 0.029}$ $15.007_{\pm 0.026}$ $14.755_{\pm 0.045}$ $14.468_{\pm 0.028}$
lending $11.184_{\pm 0.000}$ $15.130_{\pm 0.320}$ $15.112_{\pm 0.286}$ $14.957_{\pm 0.149}$ $14.916_{\pm 0.196}$ $14.962_{\pm 0.090}$
news $3.615_{\pm 0.000}$ $3.894_{\pm 0.016}$ $3.737_{\pm 0.012}$ $3.692_{\pm 0.014}$ $3.746_{\pm 0.010}$ $3.676_{\pm 0.008}$
nmes $1.931_{\pm 0.000}$ $1.202_{\pm 0.019}$ $1.014_{\pm 0.013}$ $0.958_{\pm 0.008}$ $0.952_{\pm 0.022}$ $0.780_{\pm 0.031}$
@@
adult $0.797_{\pm 0.000}$ $0.780_{\pm 0.002}$ $0.788_{\pm 0.001}$ $0.786_{\pm 0.001}$ $0.790_{\pm 0.001}$ $0.789_{\pm 0.002}$
bank $0.745_{\pm 0.002}$ $0.759_{\pm 0.004}$ $0.758_{\pm 0.006}$ $0.760_{\pm 0.005}$ $0.751_{\pm 0.005}$ $0.766_{\pm 0.005}$
churn $0.873_{\pm 0.003}$ $0.850_{\pm 0.006}$ $0.861_{\pm 0.008}$ $0.863_{\pm 0.004}$ $0.860_{\pm 0.008}$ $0.863_{\pm 0.007}$
covertype $0.817_{\pm 0.001}$ $0.791_{\pm 0.001}$ $0.745_{\pm 0.001}$ $0.761_{\pm 0.001}$ $0.768_{\pm 0.001}$ $0.767_{\pm 0.001}$
default $0.674_{\pm 0.001}$ $0.672_{\pm 0.002}$ $0.674_{\pm 0.002}$ $0.671_{\pm 0.002}$ $0.673_{\pm 0.003}$ $0.675_{\pm 0.002}$
diabetes $0.621_{\pm 0.002}$ $0.616_{\pm 0.002}$ $0.612_{\pm 0.003}$ $0.612_{\pm 0.003}$ $0.617_{\pm 0.002}$ $0.619_{\pm 0.002}$
@@
adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.909_{\pm 0.000}$ $0.909_{\pm 0.000}$ $0.909_{\pm 0.001}$ $0.909_{\pm 0.001}$
bank $0.947_{\pm 0.000}$ $0.945_{\pm 0.003}$ $0.946_{\pm 0.002}$ $0.945_{\pm 0.002}$ $0.946_{\pm 0.000}$ $0.946_{\pm 0.000}$
churn $0.964_{\pm 0.001}$ $0.959_{\pm 0.003}$ $0.961_{\pm 0.002}$ $0.962_{\pm 0.002}$ $0.960_{\pm 0.002}$ $0.957_{\pm 0.008}$
covertype $0.892_{\pm 0.000}$ $0.870_{\pm 0.000}$ $0.826_{\pm 0.001}$ $0.838_{\pm 0.001}$ $0.844_{\pm 0.001}$ $0.844_{\pm 0.000}$
default $0.768_{\pm 0.000}$ $0.763_{\pm 0.002}$ $0.764_{\pm 0.001}$ $0.764_{\pm 0.002}$ $0.764_{\pm 0.002}$ $0.764_{\pm 0.002}$
diabetes $0.693_{\pm 0.001}$ $0.675_{\pm 0.001}$ $0.664_{\pm 0.001}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.673_{\pm 0.002}$
@@
acsincome $0.804_{\pm 0.012}$ $0.868_{\pm 0.011}$ $0.808_{\pm 0.014}$ $0.820_{\pm 0.012}$ $0.800_{\pm 0.011}$ $0.807_{\pm 0.008}$
beijing $0.711_{\pm 0.001}$ $0.801_{\pm 0.006}$ $0.780_{\pm 0.005}$ $0.771_{\pm 0.005}$ $0.769_{\pm 0.006}$ $0.766_{\pm 0.003}$
lending $0.030_{\pm 0.000}$ $0.124_{\pm 0.006}$ $0.059_{\pm 0.002}$ $0.072_{\pm 0.001}$ $0.067_{\pm 0.002}$ $0.066_{\pm 0.002}$
news $1.001_{\pm 0.002}$ $0.772_{\pm 0.019}$ $0.835_{\pm 0.062}$ $0.805_{\pm 0.079}$ $0.763_{\pm 0.062}$ $0.752_{\pm 0.067}$
nmes $1.001_{\pm 0.003}$ $0.967_{\pm 0.064}$ $1.128_{\pm 0.088}$ $1.195_{\pm 0.087}$ $1.225_{\pm 0.070}$ $1.109_{\pm 0.055}$
Appendix U Training and Sampling Times Details
@@
Table 29: Sample times in seconds per 1000 samples. TabDDPM produces NaNs during training on acsincome and diabetes, and is therefore excluded for these data. CoDi is considered prohibitively expensive to train on diabetes and lending.
@@
Figures 24 and 25 show the benefit of deep generative models over SMOTE. Even though SMOTE is often praised as a simple, easy-to-use oversampling tool for tabular data, it relies on identifying nearest neighbors, making sampling very inefficient for larger datasets. As a consequence, we deem SMOTE to be infeasible to use for the acsincome and covertype datasets. The figures also illustrate the performance edge of CDTD, in particular compared to other diffusion-based models.
Figure 24: Average sample quality metrics as a function of sampling time. Diffusion-based models are indicated in orange. Only datasets for which we could retrieve results from all models are included, this excludes acsincome, covertype, diabetes, lending.
Figure 25: Average ML efficiency metrics and DCR as a function of sampling time. Diffusion-based models are indicated in orange. The dotted line indicates the test set performance of the real data. Only datasets for which we could retrieve results from all models are included, this excludes acsincome, covertype, diabetes, lending.